OpenStudy (shubhamsrg):

infinite summation : 1 + (1+2)/2! + (1+2+3)/3! ........

OpenStudy (shubhamsrg):

or in other words 1/2 * summation ( n(n+1)/n! ) from n=1 or 0 ,,to infinity..

OpenStudy (anonymous):

Something * e

OpenStudy (anonymous):

\[e^x=\sum_{n=0}^{\infty} \frac{x^n}{n!}\]take derivative\[e^x=\sum_{n=1}^{\infty} \frac{nx^{n-1}}{n!}\]multiply by x^2\[x^2e^x=\sum_{n=1}^{\infty} \frac{nx^{n+1}}{n!}\]and ...

OpenStudy (shubhamsrg):

another derivative and then put x=1 !! thanks a lot!!